OpenStudy (anonymous):

three added to 4 times a number is equal to 5 less than 6 times the number. What is the number?

OpenStudy (calculusxy):

3 + 4n = 6n - 5 solve for n. do you know how?

OpenStudy (anonymous):

no

OpenStudy (md152727):

I can help

OpenStudy (anonymous):

ok

OpenStudy (md152727):

n=4

OpenStudy (calculusxy):

bring variable and constant each to one side of the equation. like 3n = 6 (3n is to one side and 6 is to the other) same applies to here: 3 + 4n = 6n -5 3 + 4n - 4n = 6n - 4n - 5 3 = 2n - 5 3 + 5 = 2n - 5 + 5 8 = 2n 8/2 = 2n/2 4 = n

OpenStudy (anonymous):

can you explain how you got it?

OpenStudy (md152727):

3+4 *4= 28 4*6-5= 29

OpenStudy (md152727):

hmm.....

OpenStudy (md152727):

@calculusxy can you help

OpenStudy (md152727):

i got it by mind

OpenStudy (md152727):

I tried numbers

OpenStudy (calculusxy):

you can check out the work that i provided @KatyKatie

OpenStudy (anonymous):

thanks